# Rephasing invariant formulae for general parameterizations of flavor mixing matrix and exact sum rules with unitarity triangles

### v1

@@ Abstract
In this letter, we present rephasing invariant formulae $\delta^{(\alpha i)}=\arg[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}\det V}]$ for CP phases $\delta^{(\alpha i)}$ associated with nine Euler-angle-like parameterizations of a flavor mixing matrix. Here, $\alpha$ and $i$ denote the row and column carrying the trivial phases in a given parameterization. Furthermore, we show that the phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of unitarity triangles satisfy compact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}$ and $\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ where all indices are taken cyclically modulo three. These twelve relations are natural generalizations of the previous result $\delta_{\mathrm{PDG}}+\delta_{\mathrm{KM}}=\pi-\alpha+\gamma.$
@@ I Introduction
The flavor mixing of quarks and leptons has been one of the essential pieces for the development of particle physics, and it is also an intrinsically interesting subject of mathematical investigation. Among various representations of the mixing matrix Kobayashi:1973fv ; Chau:1984fp ; Hall:1993ni ; Kaur:2023ypg ; Yang:2020qsa ; Yang:2020goc ; Yang:2024ulq ; Yang:2025yst , nine Euler-angle-like parameterizations in terms of three rotation matrices and a distinct CP phase admit nine possible forms Fritzsch:1995nx ; Rasin:1997pn . However, at first glance, these nine phases seem to take chaotic values, showing no apparent pattern.
On the other hand, unitarity conditions of the mixing matrix give rise to six unitarity triangles Wu:1994di ; Xing:2009eg ; Harrison:2009bz ; He:2013rba ; He:2016dco ; Xing:2019tsn ; Harrison:2025rkp and relations between the CP phases and the unitarity triangles has also been extensively studied Koide:2004gj ; Koide:2008yu ; Frampton:2010ii ; Dueck:2010fa ; Frampton:2010uq ; Li:2010ae ; Qin:2011bq ; Zhou:2011xm ; Qin:2011ub ; Qin:2010hn ; Zhang:2012ys ; Li:2012zxa ; Zhang:2012bk . These nine angles also seem to have independent values, showing no obvious structure.
More recently, rephasing invariant formulae for the CP phases Yang:2025hex ; Yang:2025law ; Yang:2025ftl produce a simple sum rule relating the CP phases and angles $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$. This led to a discovery of an underlying order between the random CP phases and angles.
In this letter, to establish exact sum rules connecting the nine CP phases and the nine angles of unitarity triangles, we first derive rephasing invariant formulae for each CP phase in terms of elements of the mixing matrix $V$ and $\det V$. Next, we present sum rules between the phases and angles, showing that the previously known sum rule arises as a special case of these relations.
@@ II Rephasing invariant formulae for CP phases in general parameterizations of flavor mixing matrix
where $s_{\theta}=\sin\theta,c_{\theta}=\cos\theta.$ Furthermore, we define complex rotation matrices $R_{12}(\theta,\delta)$, $R_{23}(\sigma,\delta)$, and $R_{31}(\tau,\delta)$ by replacing $1\rightarrow e^{-i\delta}$ in Eq. (1). With these matrices, the mixing matrix $V$ admits nine different parameterizations Fritzsch:1997st
$P1:~V^{(33)}$ $=R_{12}(\theta)R_{23}(\sigma,\delta^{(33)})R^{-1}_{12}(\theta^{\prime})\,,~~P2:~V^{(11)}=R_{23}(\sigma)R_{12}(\theta,\delta^{(11)})R^{-1}_{23}(\sigma^{\prime})\,$, (2)
$P3:~V^{(13)}$ $=R_{23}(\sigma)R_{31}(\tau,\delta^{(13)})R_{12}(\theta)\,,~~~P4:~V^{(31)}=R_{12}(\theta)R_{31}(\tau,\delta^{(31)})R^{-1}_{23}(\sigma)\,$, (3)
$P5:~V^{(22)}$ $=R_{31}(\tau)R_{12}(\theta,\delta^{(22)})R^{-1}_{31}(\tau^{\prime})\,,~~P6:~V^{(32)}=R_{12}(\theta)R_{23}(\sigma,\delta^{(32)})R_{31}(\tau)\,$, (4)
$P7:~V^{(12)}$ $=R_{23}(\sigma)R_{12}(\theta,\delta^{(12)})R^{-1}_{31}(\tau)\,,~~P8:~V^{(21)}=R_{31}(\tau)R_{12}(\theta,\delta^{(21)})R_{23}(\sigma)\,$, (5)
@@
Since there is a duplication for the element $V_{\alpha i}^{(\alpha i)}$, there are effectively five independent conditions, and the phase $\pi$ appears either zero or two times. Finally, if the Majorana phases are omitted,
@@
In other words, the phase structure of these parameterizations $V^{(\alpha i)}$ is specified by these six conditions. Since the resulting CP phases are manifestly rephasing invariant, this construction remains valid even in the presence of Majorana phases for leptons.
Let us consider transforming a mixing matrix $V$ in an arbitrary basis into one of these parameterizations $V^{(\alpha i)}$. Suppose that unphysical phases are removed by general phase transformations:
@@
From a relation $\arg\det V=-\delta^{(\alpha i)}+\sum_{\beta,j}(\gamma_{L\beta}-\gamma_{Rj})$ obtained by taking the argument of determinant, the CP phase $\delta^{(\alpha i)}$ is found to be
$\delta^{(\alpha i)}$ $=(\gamma_{L1}+\gamma_{L2}+\gamma_{L3}-\gamma_{R1}-\gamma_{R2}-\gamma_{R3})-\arg\det V$ (11)
@@
In practice, since $V_{\alpha i}$ appears twice in the numerator, cancellations occur between the numerator and denominator:
$\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right]=\arg\left[\prod_{\beta\neq\alpha,j\neq i}{V_{\alpha j}V_{\beta i}\over V_{\alpha i}\det V}\right].$ (13)
Since six phases are required to eliminate unphysical phases $\gamma_{L\alpha}$ and $\gamma_{Ri}$, any other combinations have nontrivial phases of $V^{(\alpha i)}_{\beta j}$ explicitly.
@@
$\delta^{(11)}$ $=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right],~~\delta^{(12)}=\arg\left[{V_{11}V_{13}V_{22}V_{32}\over V_{12}\det V}\right],~~\delta^{(13)}=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right],$ (14)
$\delta^{(21)}$ $=\arg\left[{V_{22}V_{23}V_{11}V_{31}\over V_{21}\det V}\right],~~\delta^{(22)}=\arg\left[{V_{21}V_{23}V_{12}V_{32}\over V_{22}\det V}\right],~~\delta^{(23)}=\arg\left[{V_{21}V_{22}V_{13}V_{33}\over V_{23}\det V}\right],$ (15)
@@
They provide a necessary and sufficient set of “fifth-order” invariants, because only nine such invariants can be constructed using $\det V$.
By using the CKM matrix parameters from the latest UTfit results UTfit:2022hsi
$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}=0.04200\pm 0.00047\,$, (17)
@@
$\begin{pmatrix}\delta^{(11)}&\delta^{(12)}&\delta^{(13)}\\ \delta^{(21)}&\delta^{(22)}&\delta^{(23)}\\ \delta^{(31)}&\delta^{(32)}&\delta^{(33)}\end{pmatrix}=\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.16^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}$. (19)
@@
Interestingly, a sum over any row or column of the phases $\delta^{(ij)}$ yields the same value
@@
$\cos\delta^{(11)}$ $=\frac{\left(1-|V_{11}|^{2}\right)^{2}|V_{22}|^{2}-|V_{11}V_{12}V_{21}|^{2}-|V_{13}V_{31}|^{2}}{2|V_{11}V_{12}V_{13}V_{21}V_{31}|}=-0.0425\,$, (23)
$\cos\delta^{(12)}$ $=\frac{\left(1-|V_{12}|^{2}\right)^{2}|V_{23}|^{2}-|V_{12}V_{13}V_{22}|^{2}-|V_{11}V_{32}|^{2}}{2|V_{11}V_{12}V_{13}V_{22}V_{32}|}=-0.437\,$, (24)
$\cos\delta^{(13)}$ $=-\frac{\left(1-|V_{13}|^{2}\right)^{2}|V_{22}|^{2}-|V_{12}V_{23}V_{13}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{12}V_{13}V_{23}V_{33}|}=0.420\,$, (25)
$\cos\delta^{(21)}$ $=-\frac{\left(1-|V_{21}|^{2}\right)^{2}|V_{12}|^{2}-|V_{11}V_{21}V_{22}|^{2}-|V_{23}V_{31}|^{2}}{2|V_{21}V_{22}V_{23}V_{11}V_{31}|}=-0.924\,$, (26)
$\cos\delta^{(22)}$ $=\frac{\left(1-|V_{22}|^{2}\right)^{2}|V_{33}|^{2}-|V_{22}V_{23}V_{32}|^{2}-|V_{12}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{12}V_{32}|}=0.9998\,$, (27)
$\cos\delta^{(23)}$ $=-\frac{\left(1-|V_{23}|^{2}\right)^{2}|V_{32}|^{2}-|V_{22}V_{23}V_{33}|^{2}-|V_{13}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{13}V_{33}|}=-0.421\,$, (28)
$\cos\delta^{(31)}$ $=-\frac{\left(1-|V_{31}|^{2}\right)^{2}|V_{22}|^{2}-|V_{21}V_{32}V_{31}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{21}V_{31}V_{32}V_{33}|}=0.917\,$, (29)
$\cos\delta^{(32)}$ $=-\frac{\left(1-|V_{32}|^{2}\right)^{2}|V_{23}|^{2}-|V_{22}V_{32}V_{33}|^{2}-|V_{12}V_{31}|^{2}}{2|V_{31}V_{32}V_{33}V_{12}V_{22}|}=-0.917\,$, (30)
@@
Since $\delta^{(\alpha i)}$ also carries the sign of $\sin\delta^{(\alpha i)}$, it contains more information than an evaluation based on absolute values of matrix elements.
III Sum rules between CP phases in general parameterizations and unitarity triangles
It is noteworthy to revisit relations between the results and angles of unitarity triangles. The three angles of the unitarity triangle are rephasing invariants,
@@ III Sum rules between CP phases in general parameterizations and unitarity triangles
Since solving general sum rules is a difficult task, we employ a transfer-matrix approach. Defining a phase matrix $\Delta$ and a transfer matrix $T$,
@@
This represents the transfer along the columns of the matrix elements,
$\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}\,.$ (39)
Here, the indices are taken modulo three. Note that when the transfer $T^{T}$ acts from the left, the indices of elements increase. These sum rules (39) have six independent constraints, because the relation for one of the columns is trivial. These relations are numerically verified as
$\begin{pmatrix}65.16^{\circ}&92.44^{\circ}&115.91^{\circ}\\ 114.89^{\circ}&157.51^{\circ}&1.089^{\circ}\\ 93.45^{\circ}&23.54^{\circ}&156.49^{\circ}\\ \end{pmatrix}-\begin{pmatrix}115.91^{\circ}&65.16^{\circ}&92.44^{\circ}\\ 1.089^{\circ}&114.89^{\circ}&157.51^{\circ}\\ 156.49^{\circ}&93.45^{\circ}&23.54^{\circ}\\ \end{pmatrix}=\begin{pmatrix}-50.76&27.29&23.47\\ 113.80&42.62&-156.42\\ -63.04&-69.91&132.95\\ \end{pmatrix}$ (40)
@@
Similarly, we can find a transfer along rows as
@@
$\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}\,.$ (44)
@@
$\delta^{(11)}-\delta^{(13)}=\Phi_{2,2}-\Phi_{3,2}\,,~~~\pi-\delta_{\rm KM}-\delta_{\rm PDG}=\alpha-\gamma\,.$ (45)
@@
A comprehensive study of this kind of relation will be performed in the subsequent work.
The results generalize the previous sum rule to the nine CP phases in the different parameterizations and angles of unitarity triangles. Since the constraints from these sum rules provide twelve conditions for the nine phases and nine angles, six independent fundamental observables remain.

### v2

# Rephasing invariant formulae for CP phases in general parameterizations of flavor mixing matrix and exact sum rules with unitarity triangles

@@ Abstract
In this letter, we present rephasing invariant formulae $\delta^{(\alpha i)}=\arg[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}\det V}]$ for CP phases $\delta^{(\alpha i)}$ associated with nine Euler-angle-like parameterizations of a flavor mixing matrix. Here, $\alpha$ and $i$ denote the row and column carrying the trivial phases in a given parameterization. Furthermore, we show that the phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of unitarity triangles satisfy compact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha+1,i}-\Phi_{\alpha+2,i}$ and $\delta^{(\alpha+1,i)}-\delta^{(\alpha+2,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ where all indices are taken cyclically modulo three. These relations are natural generalizations of the previous result $\delta_{\mathrm{PDG}}+\delta_{\mathrm{KM}}=\pi-\alpha+\gamma.$
@@ I Introduction
The flavor mixing of quarks and leptons has been one of the essential pieces for the development of particle physics, and it is also an intrinsically interesting subject of mathematical investigation. Among various representations of the mixing matrix Kobayashi:1973fv ; Chau:1984fp ; Hall:1993ni ; Fritzsch:1995nx ; Kaur:2023ypg ; Yang:2020qsa ; Yang:2020goc ; Yang:2024ulq ; Yang:2025yst , Euler-angle-like parameterizations in terms of three rotation matrices and a distinct CP phase admit nine possible forms Rasin:1997pn . However, at first glance, these nine phases seem to take chaotic values, showing no apparent pattern.
On the other hand, unitarity conditions of the mixing matrix give rise to six unitarity triangles Wu:1994di ; Xing:2009eg ; Harrison:2009bz ; He:2013rba ; He:2016dco ; Xing:2019tsn ; Harrison:2025rkp and a lot of papers have also been studied relations between the CP phases and the triangles Koide:2004gj ; Koide:2008yu ; Frampton:2010ii ; Dueck:2010fa ; Frampton:2010uq ; Li:2010ae ; Qin:2011bq ; Zhou:2011xm ; Qin:2011ub ; Qin:2010hn ; Zhang:2012ys ; Li:2012zxa ; Zhang:2012bk . These nine angles also seem to have independent values.
More recently, rephasing invariant formulae for the CP phases Yang:2025hex ; Yang:2025law ; Yang:2025ftl produce a simple sum rule relating the CP phases and angles $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$. This led to the discovery of an underlying order between the random CP phases and angles.
In this letter, to organize relations connecting the nine CP phases and the nine angles of unitarity triangles, we first derive rephasing invariant formulae for each CP phase in terms of elements of the mixing matrix $V$ and $\det V$. Next, we present exact sum rules between the phases and angles, demonstrating that the previous sum rule is a special case of these relations.
@@ II Rephasing invariant formulae for CP phases in general parameterizations of flavor mixing matrix
where $s_{\theta}=\sin\theta,c_{\theta}=\cos\theta.$ Furthermore, complex rotation matrices $R_{12}(\theta,\delta)$, $R_{23}(\sigma,\delta)$, and $R_{31}(\tau,\delta)$ are defined by replacing $1\rightarrow e^{-i\delta}$ in Eq. (1). With these matrices, the flavor mixing matrix $V$ admits nine different parameterizations Fritzsch:1997st
$P1:~V^{(33)}$ $=R_{12}(\theta)R_{23}(\sigma,\delta^{(33)})R^{-1}_{12}(\theta^{\prime})\,,~~P2:~V^{(11)}=R_{23}(\sigma)R_{12}(\theta,\delta^{(11)})R^{-1}_{23}(\sigma^{\prime})\,$,
$P3:~V^{(13)}$ $=R_{23}(\sigma)R_{31}(\tau,\delta^{(13)})R_{12}(\theta)\,,~~~P4:~V^{(31)}=R_{12}(\theta)R_{31}(\tau,\delta^{(31)})R^{-1}_{23}(\sigma)\,$,
$P5:~V^{(22)}$ $=R_{31}(\tau)R_{12}(\theta,\delta^{(22)})R^{-1}_{31}(\tau^{\prime})\,,~~P6:~V^{(32)}=R_{12}(\theta)R_{23}(\sigma,\delta^{(32)})R_{31}(\tau)\,$,
$P7:~V^{(12)}$ $=R_{23}(\sigma)R_{12}(\theta,\delta^{(12)})R^{-1}_{31}(\tau)\,,~~P8:~V^{(21)}=R_{31}(\tau)R_{12}(\theta,\delta^{(21)})R_{23}(\sigma)\,$,
@@
Since there is a duplication for the element $V_{\alpha i}^{(\alpha i)}$, there are effectively five independent conditions, and the phase $\pi$ appears either zero or two times. As the final condition, the argument of the determinant is
@@
In other words, the phase structures of these parameterizations $V^{(\alpha i)}$ are specified by these six conditions.
Let us consider transforming the mixing matrix $V$ in an arbitrary basis into one of these parameterizations $V^{(\alpha i)}$. Suppose that general rephasing transformations remove unphysical phases as
@@
By taking arguments of the determinants, a relation $\arg\det V=-\delta^{(\alpha i)}+\sum_{\beta,j}(\gamma_{L\beta}-\gamma_{Rj})$ is obtained and the CP phase $\delta^{(\alpha i)}$ is found to be
$\delta^{(\alpha i)}$ $=(\gamma_{L1}+\gamma_{L2}+\gamma_{L3}-\gamma_{R1}-\gamma_{R2}-\gamma_{R3})-\arg\det V$
@@
In practice, since $V_{\alpha i}$ appears twice in the numerator, cancellations occur between the numerator and denominator as
$\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right]=\arg\left[{(\prod_{j\neq i}V_{\alpha j})\,(\prod_{\beta\neq\alpha}V_{\beta i})\over V_{\alpha i}\det V}\right].$ (8)
Since six phases are required to eliminate unphysical phases $\gamma_{L\alpha}$ and $\gamma_{Ri}$, any other combinations have nontrivial phases of $V^{(\alpha i)}_{\beta j}$ and are not solved for $\delta^{(\alpha i)}$ explicitly. This derivation remains valid even in the presence of Majorana phases for leptons because the resulting CP phases are manifestly rephasing invariant.
@@
$\delta^{(11)}$ $=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right],~~\delta^{(12)}=\arg\left[{V_{11}V_{13}V_{22}V_{32}\over V_{12}\det V}\right],~~\delta^{(13)}=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right],$
$\delta^{(21)}$ $=\arg\left[{V_{22}V_{23}V_{11}V_{31}\over V_{21}\det V}\right],~~\delta^{(22)}=\arg\left[{V_{21}V_{23}V_{12}V_{32}\over V_{22}\det V}\right],~~\delta^{(23)}=\arg\left[{V_{21}V_{22}V_{13}V_{33}\over V_{23}\det V}\right],$
@@
They provide a necessary and sufficient set of “irreducible fifth-order” invariants which cannot be decomposed into second- and third-orders, because only nine such invariants are constructed using $\det V$.
From the latest UTfit results of the CKM matrix parameters UTfit:2022hsi
$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}=0.04200\pm 0.00047\,$,
@@
$\begin{pmatrix}\delta^{(11)}&\delta^{(12)}&\delta^{(13)}\\ \delta^{(21)}&\delta^{(22)}&\delta^{(23)}\\ \delta^{(31)}&\delta^{(32)}&\delta^{(33)}\end{pmatrix}=\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.15^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}$. (11)
@@
Interestingly, a sum over any row or column of the phases $\delta^{(ij)}$ yields the same value of a ninth-order invariant
@@
$\cos\delta^{(11)}$ $=\frac{\left(1-|V_{11}|^{2}\right)^{2}|V_{22}|^{2}-|V_{11}V_{12}V_{21}|^{2}-|V_{13}V_{31}|^{2}}{2|V_{11}V_{12}V_{13}V_{21}V_{31}|}=-0.0425\,$,
$\cos\delta^{(12)}$ $=\frac{\left(1-|V_{12}|^{2}\right)^{2}|V_{23}|^{2}-|V_{12}V_{13}V_{22}|^{2}-|V_{11}V_{32}|^{2}}{2|V_{11}V_{12}V_{13}V_{22}V_{32}|}=-0.437\,$,
$\cos\delta^{(13)}$ $=-\frac{\left(1-|V_{13}|^{2}\right)^{2}|V_{22}|^{2}-|V_{12}V_{23}V_{13}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{12}V_{13}V_{23}V_{33}|}=0.420\,$,
$\cos\delta^{(21)}$ $=-\frac{\left(1-|V_{21}|^{2}\right)^{2}|V_{12}|^{2}-|V_{11}V_{21}V_{22}|^{2}-|V_{23}V_{31}|^{2}}{2|V_{21}V_{22}V_{23}V_{11}V_{31}|}=-0.924\,$,
$\cos\delta^{(22)}$ $=\frac{\left(1-|V_{22}|^{2}\right)^{2}|V_{33}|^{2}-|V_{22}V_{23}V_{32}|^{2}-|V_{12}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{12}V_{32}|}=0.9998\,$,
$\cos\delta^{(23)}$ $=-\frac{\left(1-|V_{23}|^{2}\right)^{2}|V_{32}|^{2}-|V_{22}V_{23}V_{33}|^{2}-|V_{13}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{13}V_{33}|}=-0.421\,$,
$\cos\delta^{(31)}$ $=-\frac{\left(1-|V_{31}|^{2}\right)^{2}|V_{22}|^{2}-|V_{21}V_{32}V_{31}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{21}V_{31}V_{32}V_{33}|}=0.917\,$,
$\cos\delta^{(32)}$ $=-\frac{\left(1-|V_{32}|^{2}\right)^{2}|V_{23}|^{2}-|V_{22}V_{32}V_{33}|^{2}-|V_{12}V_{31}|^{2}}{2|V_{31}V_{32}V_{33}V_{12}V_{22}|}=-0.917\,$,
@@
Since $\delta^{(\alpha i)}$ also carries the sign of $\sin\delta^{(\alpha i)}$, it contains more information than such an evaluation based on absolute values of matrix elements.
III Exact sum rules between CP phases in general parameterizations and unitarity triangles
It is noteworthy to reconsider relations between the results and angles of unitarity triangles. The three angles of the unitarity triangle are rephasing invariants,
@@ III Exact sum rules between CP phases in general parameterizations and unitarity triangles
Since solving general sum rules is a challenging task, we employ a transfer matrix approach. Defining a phase matrix $\Delta$ and a transfer matrix $T$,
@@
This represents the transfer along columns of $\Delta$ and rows of $\Phi$;
$\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha+1,i}-\Phi_{\alpha+2,i}\,.$ (23)
Here, the indices are taken modulo three. Note that the transfer $T^{T}$ from the left increases the indices of rows. These relations are numerically verified as
$\begin{pmatrix}65.15^{\circ}&92.44^{\circ}&115.91^{\circ}\\ 114.89^{\circ}&157.51^{\circ}&1.089^{\circ}\\ 93.45^{\circ}&23.54^{\circ}&156.49^{\circ}\\ \end{pmatrix}-\begin{pmatrix}115.91^{\circ}&65.15^{\circ}&92.44^{\circ}\\ 1.089^{\circ}&114.89^{\circ}&157.51^{\circ}\\ 156.49^{\circ}&93.45^{\circ}&23.54^{\circ}\\ \end{pmatrix}=\begin{pmatrix}-50.76&27.29&23.47\\ 113.80&42.62&-156.42\\ -63.04&-69.91&132.95\\ \end{pmatrix}$
@@
Similarly, we can find another transfer as
@@
$\delta^{(\alpha+1,i)}-\delta^{(\alpha+2,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}\,.$ (27)
@@
$\delta^{(11)}-\delta^{(13)}=\Phi_{22}-\Phi_{32}\,,~~~\pi-\delta_{\rm KM}-\delta_{\rm PDG}=\alpha-\gamma\,.$ (28)
@@
A comprehensive study of this kind of relation will be performed in the subsequent work. The results generalize the previous sum rule to the nine CP phases in the different parameterizations and angles of unitarity triangles.